In a 100 turn coil, if the flux through each turn is (t3 - 2t)mWb, the magnitude of the induced emf in the coil at a time of 4 sec is


1
Expert's answer
2021-01-12T14:50:36-0500

We can find the induced emf from the Faraday’s law of induction. It states that the emf is given by the rate of change of the magnetic flux:


"\Epsilon=-N\dfrac{d\Phi_B}{dt},"

here, "\Epsilon" is the induced electromotive force, "\Phi_B" is the magnetic flux through a single turn of coil and "N" is the number of turns of coil.

Let’s substitute "N" and "\Phi_B" into the last formula and find the magnitude of induced electromotive force:


"\Epsilon=(-100)\cdot\dfrac{d}{dt}(t^3-2t)\ mWb,""\Epsilon=(-100)\cdot(3t^2-2)\ mWb.""\Epsilon(t=4\ s)=|(-100)\cdot(3\cdot(4\ s)^2-2)\cdot10^{-3}\ Wb|=4.6\ V."

Answer:

"\Epsilon(t=4\ s)=4.6\ V."